Two long parallel wires are separated by 5.11 cm a
Answer #1

Given: seperation between wires (d) = 5.11 cm = 5.11*10-2 m; currents through wires I1 = 1.63 A and I2 = 3.57 A

Now we know that the force acting per unit length of parallel wires is given by:

F = μ0*I1*I2/2*π*d

Hence the force acting on the wire of length 3.15 m would be

Fnet = 3.15*4π*10-7*1.63*3.57/2π*5.11*10-2 = 7.17*10-5 N